OpenStudy (anonymous):

Area=(4 1/3)(14.5) I currently cannot solve this because of the fraction, what am I to do to rid of this fraction?

OpenStudy (anonymous):

Actually, the fraction isn't the problem. Change 14.5 to 14 1/2 and then change both numbers to full fractions: \[\frac{14}{3} \times \frac{29}{2}\] You should be good to go from there.

OpenStudy (anonymous):

What if in the midst, this were a semi-perimeter problem?

OpenStudy (anonymous):

Find a common denominator.

OpenStudy (anonymous):

Hint: \[\frac{14}{3} = \frac{28}{6}\]

OpenStudy (anonymous):